OpenStudy (anonymous):

4/x-40/x+6+3=0. find all real solutions

OpenStudy (anonymous):

4/x-40/x=-36/x 6+3=9 so 9=36/x 9x=36 x=4

OpenStudy (anonymous):

there are 2 solutions i believe